1. An ideal fluid flows at 15.5ms^-1 in a horizontal pipe down the line, the pipe diameter changes to 1/3 of its original diameter what is the flow speed of the fluid in this section if the original diameter of the pipe was 6m?
1
Expert's answer
2019-08-16T09:21:14-0400

Continuity equation states


"A_1v_1=A_2v_2"

So


"v_2=\frac{A_1}{A_2}v_1=\frac{\pi d_1^2/4}{\pi d_2^2/4}v_1=\frac{d_1^2}{d_2^2}v_1=\frac{d_1^2}{d_1^2/9}v_1=9v_1"

"=9\times 15.5\:\rm{m/s}=139.5\:\rm{m/s}"